OpenStudy (anonymous):

Factor. Assume that n is a natural number. 250a^3n + 16b^3n

OpenStudy (mathstudent55):

Factor a 2 out of both terms. Then each term in the parentheses is a cubic, so you have a sum of cubes.

OpenStudy (anonymous):

@mathwiz916 I got 2(5a+2b)(25a^2 -10ab+4b^2) but it says it's wrong

OpenStudy (mathstudent55):

What happened to the n's in the exponents?

OpenStudy (mathstudent55):

\(250a^{3n} + 16b^{3n} =\) \(= 2(125a^{3n} + 8b^{3n}) =\) \(=2[(5a^n)^3 + (2b^n)^3]\) \(= 2 (5a^n + 2b^n)[(5a^n)^2 - 5a^n \times 2b^n + (2b^n)^2]\) \(= 2(5a^n + 2b^n)(25a^{2n} - 10 a^nb^n + 4b^{2n} )\) If the problem did not have the n's in the exponents, you're answer would have been correct.